OpenStudy (anonymous):

Which ordered pair in the form (a, b) is a solution of this equation? 7a – 5b = 28 A. (0, 4) B. (4, 0) C. (–2, –3) D. (–3, –2)

OpenStudy (ivettef365):

try plugging the points, example first point you have (0,4) 7a - 5b = 28 7(0) - 5(4) = 28 0 - 20 = 28 -20 = 28 not good try the others

OpenStudy (anonymous):

so if i replace the numbered pairs with the a and b sooner or later i will get the answer?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

thank you so much i get it now :D

OpenStudy (ivettef365):

your welcome :)